### v1

@@ 4.1 Observations
Multi-token prediction improves algorithmic creativity significantly. In all our datasets, we observe from Fig 3 that the algorithmic creativity of the Gemma v1 (2B) model increases significantly under multi-token prediction, with nearly a  5x factor for the discovery datasets. Note that for this, we have selected the learning rate favorable towards next-token prediction; tuning for multi-token yields further gains (Fig 15).
@@
Figure 4: Multi-token diffusion training improves algorithmic creativity (top; Eq 1) on our four open-ended algorithmic tasks, and it reduces memorization on discovery tasks but not construction tasks (bottom).
Refer to captionRefer to captionRefer to caption
@@
Figure 5: Hash-conditioning significantly improves algorithmic creativity of both next- and multi-token prediction on Gemma v1 (2B) model. The labels in the X-axis denote the prefix (used during training and inference) and the temperature (used during inference).
Figure 6: Hash-conditioning improves algorithmic creativity of the GPT-2 (86M) model (but not the diffusion model): The X-axis labels denote the training and decoding procedure, while the legend indicates the type of prefix used during both.
Hash-conditioning improves algorithmic creativity for Transformers. Orthogonal to the effect of multi-token vs. next-token objectives, we point out three crucial effects that hash-conditioning has on a Transformer. First, hash-conditioning results in the highest algorithmic creativity in both the small models (Fig 6) and the larger models (Fig 5). In fact, in our larger models, the null and pause token prefix with temperature sampling (Fig 18) exhibit almost no algorithmic creativity (they are mode collapsed, see Fig 1920). In §H.2, we find that this improved creativity from hash-conditioning comes from aiding diversity, rather than by reducing memorization. Note that we do not see gains of hash-conditioning when it comes to diffusion training (Fig 6).
Second, surprisingly, with hash-conditioning, there is no need for temperature: even greedy decoding generates diverse outputs that are as good or even better than temperature in algorithmic creativity. Besides, for any fixed temperature, prefixing a hash string only improves performance over a null prefix (Fig 618). Third, increasing the hash lengths consistently boosts algorithmic creativity for both next-token and multi-token approaches (see Fig 523).
Thus, for Transformers, we propose viewing hash-conditioning as a distinct knob for diversity with more potency than temperature-scaling. This is line with Peeperkorn et al. (2024); Chen & Ding (2023); Chung et al. (2023) who find that, in realistic tasks, temperature only has a weak correlation with creativity, often inadvertently introducing incoherence.
@@ 4.2 An initial exploration of real-world summarization
For a more realistic examination of our findings, we conduct preliminary investigation of GPT models finetuned with NTP and the multi-token teacherless objectives on summarization tasks (XSUM, CNN/DailyMail). We measure the diversity of a model for any given prompt by generating $5$ different completions and computing a Self-Bleu metric Zhu et al. (2018).
Admittedly though, a summarization task is not as open-ended as we would like: a higher quality model (i.e., higher Rouge; Lin, 2004) necessarily means lower diversity. To account for this, we plot how diversity evolves over time as a function of the quality of the model; we then find in Fig 7 that for a given model quality, the larger multi-token models achieve higher diversity (albeit only by a slight amount). This increase does not hold for smaller models and is not always noticeable for CNN/DailyMail (see §I). Interestingly, teacherless training consistently shows an increase in summarization quality, measured by Rouge.
@@ 5 Discussion
5.1 Intuition about hash-conditioning
One could view the hash prefixes as a simpler alternative to varying the wordings of a prompt Li et al. (2023); Lau et al. (2024); Naik et al. (2024) or tuning a soft-prompt (Wang et al., 2024c), both of which are known to induce diversity. But why does this help? We speculatively put forth two arguments. First, is a representational one. Fixing a random seed upfront may help the model flesh out (i.e., compute the tokens of) one thought per sample, as against maintaining a running set of multiple thoughts and computing a distribution over all their tokens at each step. A similar point is made in a concurrent position paper (Jahrens & Martinetz, 2025). The second argument is specific to next-token prediction on open-ended planning tasks: fixing a random seed upfront may help the model co-ordinate multiple interlocking random decisions in advance rather than deciding them on the fly. Finally, there are also optimization aspects of how hash-conditioning works that we do not understand (see §B.1). Regardless, it remains to be seen whether hash-conditioning is useful in tasks beyond the minimal ones we design.
@@ 5.1 Intuition about hash-conditioning
Our argument is limited to learning open-ended tasks in a supervised manner. While we do not comment on how well other approaches like RL DeepSeek-AI (2025), chain-of-thought (CoT; Wei et al., 2022), and scaling test-time compute OpenAI (2024) would fare, we remark that these methods are designed to enhance the quality of a single example. It is unclear how to design them to maximize originality against a training set, and diversity over multiple responses. Furthermore, there is a profound question as to whether merely spelling out a model’s thought in token space can be an efficient way to search and maximize diversity. This may require enumerating all possible candidates by trial and error, an impossible feat when the search space is large.
@@ 6.2 Our approach to creativity
    The type of algorithmic tasks we study capture only a tiny subset of creative tasks that fall under the taxonomy in Boden (2003). There is yet another class called transformative creativity that we do not look at, and also other important taxonomies such as the Big-C/little-c creativity Csikszentmihalyi (1996). Big-C Creativity corresponds breakthroughs and world-changing ideas; what we focus on is adjacent to a class of little-c creativity tasks. Relatedly, many real-world creative tasks appear to be “out-of-distribution” in nature, which we do not capture.
@@ 7 Related Work
Directly related to us are Khona et al. (2024); Allen-Zhu & Li (2023b) who study diversity of next-token-trained models on an open-ended algorithmic task. Khona et al. (2024) consider path-connectivity on a knowledge graph. They observe that under temperature-scaling, diversity is at odds with accuracy. We show that this tradeoff can be greatly improved when we consider alternative training methods (multi-token, or hash-conditioning). Allen-Zhu & Li (2023b) empirically demonstrate that next-token predictors are able to learn a synthetic, challenging CFG, in the “infinite” data regime ($\approx 100m$ tokens). Our datasets are not CFGs, with the exception of Sibling Discovery, which can be thought of as a simple PCFG. Our negative result does not contradict theirs since what we show is a sub-optimality of NTP in a smaller data regime. Our work also extends the above works by studying limitations in much more minimal tasks that require as little as 2-hop lookahead. There are other works that study Transformers on non-open-ended graph-algorithmic tasks, discussed in §J.
@@
Generative diversity has long been a major goal, at least until the revolution in reasoning of language models, when accuracy took prominence over diversity. Much work has gone into concerns such as mode collapse (Che et al., 2017) or posterior collapse (Bowman et al., 2016) and memorization. In LLMs, regurgitation of training data has been a serious concern (Carlini et al., 2020; 2023; Nasr et al., 2023). Our results on hash-conditioning are also reminiscent of a line of work on reinforcement learning (RL) showing that adding noises to the policy model parameters enables more efficient exploration than directly adding noises to the output space Plappert et al. (2017); Fortunato et al. (2017). We defer discussion of theoretical studies of diversity and memorization §J, along with empirical studies of creativity in natural language tasks.
@@
There has been a recent emerging discussion surrounding the role of NTP as foundational piece in developing intelligent models. On the critical side, arguments have been made about the inference-time issues with auto-regression (Dziri et al., 2024; LeCun, 2024; Kääriäinen, 2006; Ross & Bagnell, 2010). Others have reported the planning and arithmetic limitations of next-token trained models (McCoy et al., 2023; Momennejad et al., 2023; Valmeekam et al., 2023a; b; c; Bachmann & Nagarajan, 2024) where the goal is accuracy, not diversity. As for diffusion as an alternative to NTP, our findings parallel that of Ye et al. (2024) who show that their variant of diffusion is able to solve the challenging path-star task of B&N’24. We provide references to more lines of multi-token prediction work in §J
There are also other Transformer failures such as the reversal curse (Allen-Zhu & Li, 2023a) or shortcut-learning (Dziri et al., 2024; Zhang et al., 2023; Liu et al., 2023; Young & You, 2022; Lai et al., 2021; Ranaldi & Zanzotto, 2023), however these are out-of-distribution failures; the sub-optimality we show is in-distribution, like in B&N’24.
@@
Most related to hash-conditioning is DeSalvo et al. (2024) who induce diversity by varying a soft-prompt learned using a reconstruction loss. Our approach requires no modification to the architecture or the loss; however, we train the whole model, which is more expensive than training only a soft-prompt generator. A concurrent position paper (Jahrens & Martinetz, 2025) conceptually suggests injecting noise with the same motivation as us. The benefits of hash-conditioning may also be related to the fact that varying the wording in a prompt is known to induce diverse outputs (Li et al., 2023; Lau et al., 2024; Naik et al., 2024). Various works also inject noise into a Transformer, in a different form from ours (e.g., inducing Gaussian noise), and for a different function such as quality, robustness (Hua et al., 2022; Jain et al., 2024) or efficiency (Wang et al., 2024c).
@@ 8 Conclusions
This work provides a minimal test-bed of tasks abstracting distinct modes of creativity. While these tasks are admittedly an extreme caricaturization of real-world tasks, they enable us to quantify otherwise elusive metrics like originality and diversity. They also enable us to control and investigate distinct parts of the current apparatus for language modeling (next-token learning and softmax-based temperature sampling) and advocate for alternatives (multi-token learning and hash-conditioning). The surprising effectiveness of hash-conditioning raises various open questions (§B.1). There are also other profound questions as to whether reasoning-enhancing methods like RL and CoT are optimal for enhancing open-ended diversity and originality (§5.2). Overall, we hope our work inspires discussion in the various directions of multi-token prediction, creativity and planning.
@@ 9 Impact Statement
10 Acknowledgements
We wish to thank Gregor Bachmann, Jacob Springer, and Sachin Goyal for extensive feedback on a draft of the paper. We also wish to thank Mike Mozer, Suhas Kotha, Clayton Sanford, Christina Baek, Yuxiao Qu, and Ziqian Zhong for valuable early discussions and pointers. The work was supported in part by Cisco, Apple, Google, OpenAI, NSF, Okawa foundation and Schmidt Sciences.
@@ B.1 Style of noise-injection
Our technique of injecting noise into the model is somewhat different from how noise is introduced in traditional VAEs (Kingma & Welling, 2014) or GANs (Goodfellow et al., 2020), and this difference is worth noticing. In traditional approaches, although the model learns a noise-output mapping, this mapping is enforced only at a distribution level i.e., the distribution of noise vectors must map to a distribution of real vectors. However, in our approach we arbitrarily enforce what noise vector goes to what real datapoint, at a pointwise level. This raises the open questions of why hash-conditioning works in the first place — surprisingly, without breaking optimization or generalization — and whether there is a way to enforce it at distribution-level, and whether that can provide even greater improvements.
@@ B.3 Examples of Triangle Discovery
Improved algorithmic creativity is not due to some form of capacity control. While §2.6 argues that multi-token prediction should help creativity by providing critical lookahead capabilities, it is also possible that it simply acts as a form of capacity control that prevents memorization. We rule this out in Fig 8: even as memorization computed on unseen hash strings is controlled, the multi-token model perfectly reproduces the training data on seen hash strings. We term this hash-memorization. An exact equivalence of this phenomenon was noticed in GANs in Nagarajan et al. (2018), where the generator can be trained on specific latent vectors to memorize the mapping on those, and yet produce fresh samples outside of those latent vectors.
@@
Figure 8: Even if multi-token prediction reduces memorization (on unseen hash strings), it has enough capacity to memorize training data on the seen hash-strings (denoted by hash-memorization). Note that the best algorithmic creativity for NTP and MTP are achieved at step 10k and 40k, respectively, which are the checkpoints we used to report metrics in Fig 4.
@@ Appendix D Further experimental details
We represent the main tokens in our tasks with integers (ranging upwards of 00 to as many distinct integers as required). In the hash-conditioning setting, we use hash strings of default length $10$, using randomly sampled uppercase characters from the English alphabet. In all datasets, we space-separate the vertices in a string, and comma-separate the edges.
@@
Details for GPT-2 (86M) model.  We use GPT-2 (small) with 86M non-embedding parameters when we are comparing Transformers with diffusion models. We train these models with a learning rate of $10^{-4}$ and a batch size of $64$, to convergence in terms of the algorithmic creativity. We provide sensitivity analysis of learning rate in §F.
Details for SEDD (90M) model.  We use SEDD’s “absorb” variant, which begins denoising with a fully masked sequence and iteratively refines tokens over 128 denoising steps. This variant achieves the best language modeling performance in the original paper. Same as GPT-2 (86M), we train these models with a learning rate of $10^{-4}$ and a batch size of $64$, to convergence in terms of algorithmic creativity.  [^3] We use the codebase of Lou et al. (2023) at https://github.com/louaaron/Score-Entropy-Discrete-Diffusion. We provide sensitive analysis of learning rate in §F.

### v2

@@ 4.1 Observations
Multi-token prediction improves algorithmic creativity significantly. In all our datasets, the algorithmic creativity of the Gemma v1 (2B) model increases significantly under multi-token prediction ( Fig 3), with nearly a  5x factor for the discovery datasets. Note that for this, we have selected the learning rate favorable towards next-token prediction; tuning for multi-token yields further gains (Fig 15).
Refer to captionRefer to captionRefer to caption
Figure 4: Multi-token diffusion training improves algorithmic creativity (top; Eq 1) on our four open-ended tasks, and reduces memorization on discovery tasks but not construction tasks (bottom).
@@
Figure 5: Seed-conditioning significantly improves algorithmic creativity of both next- and multi-token prediction on Gemma v1 (2B) model. The X-axis labels denote the prefix (at training and inference) and the temperature (at inference).
Figure 6: Seed-conditioning improves algorithmic creativity of the GPT-2 (86M) model: All models are trained with NTP. The color denotes temperature, while the X-axis, the prefix. Remarkably, seed-conditioning achieves high creativity even with greedy decoding. However, note that for the discovery datasets, top-K sampling must be enabled to see a positive gap between seed- and null- settings.
Seed-conditioning improves algorithmic creativity for Transformers. Orthogonal to the effect of multi-token vs. next-token objectives, we point out the effect of seed-conditioning a Transformer. First, and most remarkably, with seed-conditioning, there is no need for temperature: even deterministic generation via greedy decoding generates diverse outputs that have highly non-trivial creativity values (Fig 5, Fig 6). This means that, even though the seeds we provide are arbitrarily paired with the training datapoint, the models have learned to transform the arbitrary noise into meaningful randomness. We also find that increasing the seed lengths consistently boosts algorithmic creativity for both next-token and multi-token approaches (Fig 523). Thus, for Transformers, we propose viewing seed-conditioning as a knob for diversity distinct from temperature sampling.
Next, in comparison to temperature sampling, in our settings, seed-conditioning results in the highest algorithmic creativity in both the small models (Fig 6) and the larger models (Fig 5). In fact, in our larger models, the null and pause token prefix with temperature sampling (Fig 18) exhibit almost no algorithmic creativity (they are mode collapsed, see Fig 1920). Besides, for any fixed temperature, prefixing a seed only improves performance over a null prefix (Fig 618).
In §H.2, we find that the improved creativity from seed-conditioning comes from improved diversity, rather than from reduced memorization (unlike with multi-token training). In some of our settings, this positive gap between seed-conditioning and temperature sampling is attributable to the fact that top-$K$ sampling is enabled by default as is standard (specifically, for GPT-2 (86M) on Sibling Discovery and Triangle Discovery where the vocabulary size is larger than $K$). Intuitively, top-$K$ sampling restricts the degree of freedom in output randomness, allowing the seed to introduce new degrees of freedom. However, in all other settings, this explanation does not hold. Regardless all these behaviors of seed-conditioning require further study as they are non-trivial, especially given that the seeds are arbitrarily chosen in relationship to the output. As a starting point, we provide some preliminary hypotheses for when and why seed-conditioning may help Transformers in §5.1. Note that we do not see gains of seed-conditioning when it comes to diffusion training (§ F.3).
@@ 4.2 An initial exploration of real-world summarization
For a more realistic examination of our findings, we conduct preliminary investigation of GPT models finetuned with NTP and the multi-token teacherless objectives on summarization tasks (XSUM, CNN/DailyMail). We measure the diversity of a model for any given prompt by generating $5$ different completions and computing a Self-Bleu metric (Zhu et al., 2018). Admittedly though, a summarization task is not as open-ended as we would like: a higher quality model (i.e., higher Rouge; Lin, 2004) necessarily means lower diversity. To account for this, we plot how diversity evolves over time as a function of the quality of the model; we then find in Fig 7 that for a given model quality, the larger multi-token models achieve higher diversity (albeit only by a slight amount). This increase does not hold for smaller models and is not always noticeable for CNN/DailyMail (see §I). Interestingly, teacherless training consistently shows an increase in summarization quality, measured by Rouge.
@@ 5 Discussion
5.1 Intuition about seed-conditioning
Seed-conditioning produces non-trivial algorithmic creativity even with greedy decoding. Perhaps, one could view the seed prefixes as a simpler alternative to varying the wordings of a prompt Li et al. (2023); Lau et al. (2024); Naik et al. (2024) or tuning a soft-prompt (Wang et al., 2024c), both of which are known to induce diversity. However, it is unexpected that the seeds provide any meaningful randomness at all, given they are arbitrarily picked with no semantic relationship to the training point. We do not understand this presently.
Another pertinent question is to why seed-conditioning works better than temperature sampling in the settings it does. To this, we offer three potential explanations. The simplest explanation is that in our GPT-2 (86M) settings, top-K sampling is enabled by default. This restricts the degree of freedom in the support induced by the output-layer randomness; thus, providing different seeds can help vary this support and produce more varying samples. However, this does not explain the gap we see in the other datasets where the output vocabulary is smaller than $K$ ($K=50$), or in Gemma v1 (2B) where we do not use top-K sampling.
Thus, we speculatively put forth two more arguments. First, is a representational one. Fixing a random seed upfront may help the model flesh out (i.e., compute the tokens of) one thought per sample, as against maintaining a running set of multiple thoughts and computing a distribution over all their tokens at each step. A similar point is made in a concurrent position paper (Jahrens & Martinetz, 2025). The second argument is specific to next-token prediction on open-ended planning tasks: fixing a random seed upfront may help the model co-ordinate multiple interlocking random decisions in advance rather than deciding them on the fly.
Finally, there are also optimization aspects of how seed-conditioning works that we do not understand (see §B.1). Regardless, it remains to be seen whether seed-conditioning is useful in tasks beyond the minimal ones we design.
@@ 5.1 Intuition about seed-conditioning
Our argument is limited to learning open-ended tasks in a supervised manner. While we do not comment on how well other approaches like RL (DeepSeek-AI, 2025), chain-of-thought (CoT; Wei et al., 2022), and scaling test-time compute (OpenAI, 2024) would fare, we remark that these methods are designed to enhance the quality of a single example. It is unclear how to design them to maximize originality against a training set, and diversity over multiple responses. Furthermore, there is a profound question as to whether merely spelling out a model’s thought in token space can be an efficient way to search and maximize diversity. This may require enumerating all possible candidates by trial and error, an impossible feat when the search space is large.
@@ 6.1 Limitations of our experimental conclusions
  6. 6.
    While seed-conditioning appears to be a competitive alternative to temperature sampling, there are caveats here:
    1. (a)
      In our GPT-2 (86M) experiments for Sibling Discovery and Triangle Discovery, the gap between seed-conditioning and temperature-sampling is contingent on top-K sampling (which is turned on by default in the library as it is standard practice in real-world applications). Although this does not seem to be a requirement in our other settings (and is immaterial for the other datasets which have a smaller vocabulary), it is not clear how crucial this factor would be for demonstrating the same gap in real settings.
    2. (b)
      It is not clear how well this idea may generalize to realistic data.
    3. (c)
      Unfortunately, seed-conditioning requires special training of the model, which is not as appealing as the far simpler temperature sampling.
@@ 6.2 Our approach to creativity
    The type of algorithmic tasks we study capture only a tiny subset of creative tasks that fall under the taxonomy in Boden (2003). There is yet another class called transformative creativity that we do not look at, and also other important taxonomies such as the Big-C/little-c creativity (Csikszentmihalyi, 1996). Big-C Creativity corresponds breakthroughs and world-changing ideas; what we focus on is adjacent to a class of little-c creativity tasks. Relatedly, many real-world creative tasks appear to be “out-of-distribution” in nature, which we do not capture.
@@ 7 Related Work
Directly related to us are Khona et al. (2024); Allen-Zhu & Li (2023b) who study diversity of next-token-trained models on an open-ended algorithmic task. Khona et al. (2024) demonstrate a diversity-accuracy tradeoff under temperature-scaling for path-connectivity on on a knowledge graph. We show how this tradeoff can be improved under alternative training methods. Allen-Zhu & Li (2023b) empirically demonstrate that next-token predictors are able to learn a synthetic, challenging CFG, in the “infinite” data regime ($\approx 100m$ tokens). Our datasets are not CFGs, with the exception of Sibling Discovery. Our negative result does not contradict theirs since what we show is a sub-optimality of NTP in a smaller data regime. Our work also extends the above works by studying limitations in much more minimal tasks that require as little as 2-hop lookahead. There are other works that study Transformers on non-open-ended graph-algorithmic tasks, discussed in §J.
@@
Generative diversity has long been a major goal, at least until the revolution in (deterministic) reasoning of language models. Much work has gone into concerns such as mode collapse (Che et al., 2017) or posterior collapse (Bowman et al., 2016) and memorization (Carlini et al., 2020, 2023; Nasr et al., 2023). Temperature sampling has been known to have weak correlation with creativity, often inadvertently introducing incoherence (Peeperkorn et al., 2024; Chen & Ding, 2023; Chung et al., 2023). Our results on seed-conditioning are also reminiscent of a line of work on reinforcement learning (RL) showing that adding noises to the policy model parameters enables more efficient exploration than directly adding noises to the output space (Plappert et al., 2017; Fortunato et al., 2017). We defer discussion of theoretical studies of diversity and memorization and with empirical studies on natural-language creativity in §J.
@@
There has been a recent emerging discussion surrounding the role of NTP as foundational piece in developing intelligent models. On the critical side, arguments have been made about the inference-time issues with auto-regression (Dziri et al., 2024; LeCun, 2024; Kääriäinen, 2006; Ross & Bagnell, 2010). Others have reported the planning and arithmetic limitations of next-token trained models (McCoy et al., 2023; Momennejad et al., 2023; Valmeekam et al., 2023a, b, c; Bachmann & Nagarajan, 2024) where the goal is accuracy, not diversity. As for diffusion as an alternative to NTP, our findings parallel that of Ye et al. (2024) who show that their variant of diffusion is able to solve the challenging path-star task of B&N’24. We provide references to more lines of multi-token prediction work in §J. There are also other Transformer failures such as the reversal curse (Allen-Zhu & Li, 2023a) or shortcut-learning (Dziri et al., 2024; Zhang et al., 2023; Liu et al., 2023; Young & You, 2022; Lai et al., 2021; Ranaldi & Zanzotto, 2023), however these are out-of-distribution failures; the sub-optimality we show is in-distribution, like in B&N’24.
@@
Most related to seed-conditioning is DeSalvo et al. (2024) who induce diversity by varying a soft-prompt learned using a reconstruction loss. Our approach requires no modification to the architecture or the loss; however, we train the whole model, which is more expensive. A concurrent position paper (Jahrens & Martinetz, 2025) conceptually suggests injecting noise with the same motivation as us. The benefits of seed-conditioning may also be related to the fact that varying the wording in a prompt is known to induce diverse outputs (Li et al., 2023; Lau et al., 2024; Naik et al., 2024). Various works also inject noise into a Transformer, in a different form (e.g., inducing Gaussian noise), and for a different function such as quality, robustness (Hua et al., 2022; Jain et al., 2024) or efficiency (Wang et al., 2024c).
@@ 8 Conclusions
This work provides a minimal test-bed of tasks abstracting distinct modes of creativity. While these tasks are admittedly an extreme caricaturization of real-world tasks, they enable us to quantify otherwise elusive metrics like originality and diversity. They also enable us to control and investigate distinct parts of the current apparatus for language modeling (next-token learning and softmax-based temperature sampling) and advocate for alternatives (multi-token learning and seed-conditioning). The surprising effectiveness of seed-conditioning raises various open questions (§B.1). There are also other profound questions as to whether reasoning-enhancing methods like RL and CoT are optimal for enhancing open-ended diversity and originality (§5.2). Overall, we hope our work inspires discussion in the various directions of multi-token prediction, creativity and planning.
@@ B.1 Style of noise-injection
Our technique of injecting noise into the model is somewhat different from how noise is introduced in traditional VAEs (Kingma & Welling, 2014) or GANs (Goodfellow et al., 2020), and this difference is worth noticing. In traditional approaches, although the model learns a noise-output mapping, this mapping is enforced only at a distribution level i.e., the distribution of noise vectors must map to a distribution of real vectors. However, in our approach we arbitrarily enforce what noise vector goes to what real datapoint, at a pointwise level. This raises the open questions of why seed-conditioning works in the first place — surprisingly, without breaking optimization or generalization — and whether there is a way to enforce it at distribution-level, and whether that can provide even greater improvements.
@@ B.3 Examples of Triangle Discovery
Improved algorithmic creativity is not due to some form of capacity control. While §2.6 argues that multi-token prediction should help creativity by providing critical lookahead capabilities, it is also possible that it simply acts as a form of capacity control that prevents memorization. We rule this out in Fig 8: even as memorization computed on unseen seeds is controlled, the multi-token model perfectly reproduces the training data on seen seeds. We term this seed-memorization. An exact equivalence of this phenomenon was noticed in GANs in Nagarajan et al. (2018), where the generator can be trained on specific latent vectors to memorize the mapping on those, and yet produce fresh samples outside of those latent vectors.
@@
Figure 8: Even if multi-token prediction reduces memorization (on unseen seeds), it has enough capacity to memorize training data on the seen seeds (denoted by seed-memorization). Note that the best algorithmic creativity for NTP and MTP are achieved at step 10k and 40k, respectively, which are the checkpoints we used to report metrics in Fig 4.
@@ Appendix D Further experimental details
We represent the main tokens in our tasks with integers (ranging upwards of 00 to as many distinct integers as required). In the seed-conditioning setting, we use seeds of default length $10$, using randomly sampled uppercase characters from the English alphabet. In all datasets, we space-separate the vertices in a string, and comma-separate the edges.
@@
Details for GPT-2 (86M) model.  We use GPT-2 (small) with 86M non-embedding parameters when we are comparing Transformers with diffusion models. We train these models with a learning rate of $10^{-4}$ and a batch size of $64$, to convergence in terms of the algorithmic creativity. In temperature sampling, we use nucleus sampling with a parameter of $0.95$, although we observed this does not affect the results qualitatively. We used top-k sampling with a parameter of $50$. We also explore temperatures above $1.0$ for these models since temperatures low than $1.0$ only did worse. We provide sensitivity analysis of learning rate in §F.
Details for SEDD (90M) model.  We use SEDD’s “absorb” variant, which begins denoising with a fully masked sequence and iteratively refines tokens over 128 denoising steps. This variant achieves the best language modeling performance in the original paper. Same as GPT-2 (86M), we train these models with a learning rate of $10^{-4}$ and a batch size of $64$, to convergence in terms of algorithmic creativity.  [^1] We use the codebase of Lou et al. (2023) at https://github.com/louaaron/Score-Entropy-Discrete-Diffusion. We provide sensitive analysis of learning rate in §F. For sampling, we used the default function parameters, which turned out to involve top-$K$ sampling with $K=50$.